# WIMP/FIMP dark matter and primordial black holes with memory burden effect

### v3

@@ 4.1 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. The FO of WIMPs occurs at [2]
$x_{\rm FO}=\frac{m_{\chi}}{T_{\rm FO}}\simeq 20,$ (34)
where $T_{\rm FO}$ is the FO temperature. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time as $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is
@@
The potential impact of PBH evaporation without the memory burden effect on the production of DM particles via thermal processes has been deeply investigated. In the context of the FO or FI mechanism, the authors in Ref. [50] considered fermionic DM interacting with SM particles via a spin-1 mediator. They focused on a case where the mediator is heavier than the DM but has a smaller mass than the reheating temperature.
In the case of the FO mechanism, the authors confirmed that DM particles produced from PBH evaporation thermalize with SM particles if the PBHs evaporate before FO. Therefore, PBHs do not affect the FO mechanism dynamics or contribute to the final relic abundance. Conversely, when PBH evaporation occurs after FO, two scenarios are possible.
In the first scenario, PBHs may dominate the universe’s energy density before evaporating. With PBHs dominating the universe’s energy density, PBHs act as a DM source by continuously injecting entropy into the visible sector, possibly affecting thermal production. If FO occurs during or before a PBH-dominated era, then PBH evaporation can alter standard FO predictions during PBH domination. However, unless particles from PBHs interact efficiently with thermal DM remnants and destroy FO predictions, PBHs cannot simultaneously affect FO production dynamics and contribute significantly to the relic density if they evaporate after FO. This is due to the presence of the mediator, which provides additional annihilation channels for DM and causes resonant interactions between DM and SM particles. The interaction between the evaporated DM particles and the preexisting relic abundance or between the evaporated DM particles and SM particles must be efficient for PBHs to dominate the universe’s energy density before evaporating. In this case, the Boltzmann equation should be solved at the level of phase-space distributions because of the nontrivial phase-space distribution of the evaporation products [50].
In the second scenario, PBHs never dominate the energy density. In this case, the FO mechanism and PBH evaporation can both contribute without affecting each other.
In this study, we focus on a case where the effect of the thermal production of WIMPs/FIMPs considerably exceeds that of PBH particle production via Hawking radiation. In addition, we assume that no early-matter-dominated era exists and PBHs never dominate the energy density in the history of the universe (we focus on a scenario where $\beta<\beta_{\rm c}$ holds). Therefore, DM particles from PBH evaporation are expected not to interact with thermally produced WIMPs.
To ensure that thermalization does not occur, we evaluate the interaction rates between particles from PBHs and thermally produced particles and compare them with the Hubble scale, as done in Ref. [50]. However, the current study does not examine the details of this problem. We take the range of $q$ at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$.
@@
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$. The present-day density of DM is
@@
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}$
$\quad+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
@@
We study the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$. In the following numerical calculations, we assume $k=1$ as the simplest case. We regard $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV as the benchmark. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$.
@@
at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ from Eq.(39). Thus, a small $m_{\chi}$ with a large $q$ increases $\beta$ and eliminates it via $\beta<\beta_{\rm c}$. In principle, the WDM constraint may also contribute to the lower limit of $m_{\chi}$. However, the WDM constraint is weaker than the $\beta<\beta_{\rm c}$ constraint.
@@ 4.2 FIMPs and PBHs
In this study, PBHs never dominate the energy density in the history of the universe. DM particles from PBH evaporation are expected not to interact with thermally produced FIMPs, as in the WIMP case. To ensure that thermalization does not occur in the FI case, we compare the interaction rates between particles from PBHs and thermally produced particles with the Hubble scale [50]. However, instead of this estimation, we consider the range of parameters $q$ at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$.
@@
at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ in the FIMP case. Therefore, $\beta$ is independent of $m_{\chi}$ and the relation $\beta<\beta_{\rm c}$ does not constrain $m_{\chi}$. We obtain $21.2\leq m_{\chi}^{\rm min}[{\rm GeV}]\leq 40.7$ at $0.9\geq q\geq 0$ with $M_{\rm in}=1\times 10^{6}$ from the WDM constraint.
@@ 5 Summary
The lifetime of PBHs is extended by the memory burden effect. Hence, light PBHs may survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs/FIMPs, WIMPs/FIMPs produced via the Hawking radiation of PBHs, and PBHs that survived via the memory burden effect. We can omit the contribution of thermally produced WIMPs/FIMPs to DM relic abundance, obtaining the case studied by [25], if the effect of the thermal production of WIMPs/FIMPs is much smaller than that of PBH particle production via Hawking radiation. Thus, we focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of PBH particle production via Hawking radiation.
The relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect, $q\lesssim\mathcal{O}(10^{-3})$, is only allowed at $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$ for DM consisting of thermal WIMPs, WIMPs from PBHs, and survived PBHs. A similar result is obtained if we consider FIMPs instead of WIMPs as the DM components.
We also find that the WDM constraint–the lower limit of the DM mass, which is related to structure formation in the universe–decreases with the memory burden effect of PBHs. The WDM constraint is more effective for FIMPs than for WIMPs.

### v4

@@ 4 WIMP/FIMP DM and PBHs
4.1 Thermalization of PBH-produced dark matter
In general, DM particles injected non-thermally from PBHs may thermalize with the Standard Model plasma if their interaction rate exceeds the Hubble expansion rate, as discussed for example in Ref.[50].
Before proceeding, we clarify under which conditions the DM particles produced via PBH evaporation do not affect the standard FO (or FI) mechanism. In this work, we focus on a restricted parameter region not only characterized by $\beta<\beta_{c}$, such that PBHs never dominate the energy density of the Universe, but also by
$\Omega_{\rm FO(FI)}\gg\Omega_{\rm ev},$ (53)
so that the contribution of PBH evaporation to the DM abundance is always subdominant.
We now derive a condition under which DM particles produced via PBH evaporation do not thermalize with the Standard Model plasma. Since the FO condition is defined by $\Gamma_{\rm FO}\simeq H$, it follows immediately that if
$\Gamma_{\rm ev}\ll H$ (54)
is satisfied then the DM particles produced via PBH evaporation do not enter thermal equilibrium, and the relic abundance obtained from the FO mechanism remains unaltered. The interaction rate can be estimated as
$\Gamma_{\rm ev}\sim n_{\chi}^{\rm ev}\langle\sigma v\rangle,$ (55)
where $n_{\chi}^{\rm ev}$ denotes the number density of DM particles injected by PBH evaporation and $\langle\sigma v\rangle$ is the thermally averaged interaction cross section for DM interaction with the Standard Model.
In the case where PBHs never dominate the energy density of the Universe ($\beta<\beta_{c}$), PBH evaporation takes place during a radiation-dominated era. The number density of DM particles produced by PBHs can then be estimated as
$n_{\chi}^{\rm ev}(T)\sim\beta\,\frac{\rho_{r}(T)}{M_{\rm in}}\,N_{\chi},$ (56)
where $\rho_{r}(T)\sim T^{4}$. Using the Hubble rate in the radiation-dominated era, $H(T)\sim T^{2}/M_{\rm P}$, the condition $\Gamma_{\rm ev}\ll H$ can be written as
$\beta N_{\chi}\langle\sigma v\rangle\frac{M_{\rm P}}{M_{\rm in}}\frac{m_{\chi}^{2}}{x_{\rm FO}^{2}}\ll 1,$ (57)
where we think the FO of WIMPs occurs at [2]
$x_{\rm FO}=\frac{m_{\chi}}{T_{\rm FO}}\simeq 20,$ (58)
where $T_{\rm FO}$ is the FO temperature. When the condition in Eq.(57) is satisfied, the interaction rate of PBH-produced DM particles always remains well below the Hubble expansion rate, and thermalization does not occur. In this case, the relic abundance generated by the FO (or FI) mechanism is not modified by PBH evaporation in the parameter space considered in this work.
@@ 4.2 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time as $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is
@@
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$.
Figure 1: Dependence of the threshold $\beta_{\rm nt}$, which characterizes the condition under which particle DM produced from PBHs does not thermalize, on the coupling strength $\alpha_{\rm WIMP}$. The upper three curves correspond to the case with an initial PBH mass of $10^{6}\,\mathrm{g}$, while the lower three curves correspond to $10^{8}\,\mathrm{g}$. For a fixed PBH mass, the differences among the three curves reflect different choices of the parameter $q$.
First, we assume $k=1$ as the simplest case in the following numerical calculations. We regard $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV as the benchmark. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$.
From Eq.(57), we obtain the non-themalization condition
$\beta N_{\chi}\alpha_{\rm WIMP}^{2}\frac{M_{\rm P}}{M_{\rm in}}\frac{1}{x_{\rm FO}^{2}}\ll 1,$ (63)
for WIMPs. This inequality shows that non-thermalization of PBH-produced WIMPs can be guarantee for sufficiently small initial PBH density, small interactions, or large initial PBH mass. Also the non-thermalization threshold depends on WIMP mass, initial mass of PBH and/or, strength of the memory burden effect ($N_{\chi}$ depends on $m_{\chi}$, $M_{\rm in}$ and/or, $q$). In this study, we focus the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ (the condition of Eq.(57) as well as Eq.(63) are satisfied), and PBH-produced WIMPs do not enter thermal equilibrium. The scenario in which PBH-produced WIMPs enter thermal equilibrium lies beyond the scope of the present study. For our purposes, we require not only $\beta<\beta_{c}$ but also
$\beta\ll\beta_{\rm nt}=\frac{x_{\rm FO}^{2}}{N_{\chi}\alpha_{\rm WIMP}^{2}}\frac{M_{\rm in}}{M_{\rm P}},$ (64)
to determine the initial density of PBHs, where $\beta_{\rm nt}$ denotes the threshold for non-thermalization of PBH-origin WIMPs.
Figure  1 illustrates the dependence of the threshold $\beta_{\rm nt}$, which characterizes the condition under which particle DM produced from PBHs does not thermalize, on the coupling strength $\alpha_{\rm WIMP}$. In this figure, the DM mass is fixed to $700~\mathrm{GeV}$. We have checked that varying the DM mass by $\mathcal{O}(50\%)$ does not significantly modify the value of $\beta_{\rm nt}$. The upper three curves correspond to the case with an initial PBH mass of $10^{6}\,\mathrm{g}$, while the lower three curves correspond to $10^{8}\,\mathrm{g}$. For a fixed PBH mass, the differences among the three curves reflect different choices of the parameter $q$.
We find that, as the memory burden effect becomes stronger (i.e., as $q$ increases), the threshold $\beta_{\rm nt}$ for PBH-produced WIMPs to remain non-thermalized also increases. This is because a larger memory burden effect leads to a slower evaporation of PBHs, which suppresses the total amount of WIMPs emitted even if a larger number of PBHs were formed in the early Universe.
From Fig.1, if we adopt the range $\alpha_{\rm WIMP}=0.01\sim 0.5$ as a representative interval for phenomenologically viable WIMP couplings, then for an initial PBH abundance $\beta\ll 10^{-9}$, the condition $\beta\ll\beta_{\rm nt}$ is expected to hold over the PBH mass range and the range of the memory-burden effect considered in this work. As discussed below, the value of $\beta$ required to reproduce the observed DM abundance $\Omega_{\rm DM}h^{2}\simeq 0.12$ under our setup, $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$, lies well within this regime. Accordingly, the present analysis focuses on the parameter region in which WIMP originating from PBHs is not expected to thermalize.
Before progress this subsection, we clarify why a numerical solution of the coupled Boltzmann equations is not required in the present analysis. In general, such a treatment becomes necessary only if DM particles injected by PBH evaporation can enter thermal equilibrium with the Standard Model plasma, which requires the interaction rate to satisfy $\Gamma_{\rm ev}\gtrsim H$. Instead of exploring this boundary region, we deliberately restrict ourselves to a parameter space that satisfies a sufficient condition for non-thermalization, $\Gamma_{\rm ev}\ll H$, throughout the cosmological history relevant for FO (or FI).
Using Eq. (63), this condition can be expressed as $\beta\ll\beta_{\rm nt}$, where $\beta_{\rm nt}$ denotes the threshold initial PBH fraction above which thermalization may occur. In all benchmark scenarios considered in this work, the hierarchy $\beta/\beta_{\rm nt}\lesssim 10^{-10}$ is realized, placing the system far away from the thermalization boundary. As a consequence, DM particles emitted by PBHs never reach thermal equilibrium with the thermal bath, and the relic abundance generated by the standard FO (or FI) mechanism remains unchanged.
Solving the coupled Boltzmann equations would therefore not modify the final relic abundance within the parameter space considered here. A numerical Boltzmann analysis becomes relevant only in scenarios where PBH evaporation injects a sizable DM population close to thermal equilibrium, as studied for instance in Ref. [50], which are explicitly excluded in the present work.
The present-day density of DM is
@@
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
@@
at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ from Eq.(65). Thus, a small $m_{\chi}$ with a large $q$ increases $\beta$ and eliminates it via $\beta<\beta_{\rm c}$.
@@ 4.3 FIMPs and PBHs
In this study, PBHs never dominate the energy density in the history of the universe. DM particles from PBH evaporation are expected not to interact with thermally produced FIMPs, as in the WIMP case. We consider the range of parameters $q$ at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$.
@@
at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ in the FIMP case. Therefore, $\beta$ is independent of $m_{\chi}$ and the relation $\beta<\beta_{\rm c}$ does not constrain $m_{\chi}$.
@@
We would like to comment about gravitational production of DM from SM particles through graviton exchange. In addition to the thermal and PBH-induced production mechanisms discussed above, DM can in principle also be produced via FI by purely gravitational interactions through graviton exchange. This mechanism has been studied in detail in Ref. [114]. According to Ref. [114], the thermally averaged cross section scales as ($n=2$ case) $\langle\sigma v\rangle\simeq T^{2}/M_{\rm P}^{4}$. The production rate is therefore $R(T)\sim n_{\rm SM}^{2}\langle\sigma v\rangle\simeq T^{8}/M_{\rm P}^{4}$. During radiation domination, $H\sim T^{2}/M_{\rm P}$ and $s\sim T^{3}$, which gives $dY_{\chi}/dT\simeq T^{2}/M_{\rm P}^{3}$. Integrating up to the reheating temperature yields $Y_{\chi}\simeq T_{\rm RH}^{3}/M_{\rm P}^{3}$ and hence, the resulting relic abundance scales as
$\Omega_{\rm FI}^{\rm grav}h^{2}\simeq\frac{m_{\chi}\,T_{\rm RH}^{3}}{M_{\rm P}^{4}}.$ (71)
In this study, we adopt a simple FIMP scenario, in which the FIMP relic abundance $\Omega_{\rm FI}h^{2}$ is given by Eq. (68). In the regime $\Omega_{\rm FI}^{\rm grav}h^{2}\ll\Omega_{\rm FI}h^{2}$, namely
$m_{\chi}\ll\left(\frac{1.48\times 10^{31}{\rm GeV}}{T_{\rm RH}}\right)^{3}\alpha_{\rm FIMP}^{2}~{\rm GeV},$ (72)
the contribution from graviton-mediated DM production is expected to be subdominant. Although the reheating temperature is model dependent, it is roughly estimated to lie in the range from a few GeV to $10^{15}$ GeV. Under these considerations, at least for $m_{\chi}\ll 3.2\times 10^{48}\alpha^{2}_{\rm FIMP}$ GeV, the graviton-mediated contribution is not expected to play a significant role in the present analysis. This condition is satisfied in the parameter region considered in this study.
4.4 $k$-dependence
So far, the analysis has not explicitly explored different values of the parameter $k$. Instead of varying $k$, we have focused on changes in the parameter $q$. In the following, we discuss how varying $k$ may affect the parameter space and potentially modify the results.
The parameter $k$ enters the cosmological evolution primarily through the modification of the PBH evaporation timescale and temperature, $t_{\rm ev}(k)$ and $T_{\rm ev}(k)$, as discussed in Sec. 2. At the level of parametric dependence, increasing $k$ delays the evaporation and lowers $T_{\rm ev}$, thereby increasing the typical momentum of DM particles emitted at evaporation relative to the plasma temperature.
This effect propagates into the relic abundance and the WDM constraint mainly through $T_{\rm ev}$. For instance, the free-streaming length and the corresponding lower bound on the DM mass scale approximately as $m_{\chi}^{\rm WDM}\propto T_{\rm ev}^{-1}(k)$, implying that a larger value of $k$ strengthens the WDM constraint. Apart from this indirect dependence, no additional qualitative modification of the production mechanism arises from varying $k$ within the range of interest. Quantitatively, increasing $k$ shifts the WDM excluded region toward larger DM masses through the reduction of $T_{\rm ev}$, while leaving the structure of the allowed parameter space unchanged.
In the present work, we therefore fix $k$ and instead vary the parameter $q$, which directly controls the strength of the memory burden effect, in order to efficiently illustrate its phenomenological impact. A systematic scan over $k$ would lead to a quantitative reshaping of the parameter space but would not alter the qualitative conclusions of our analysis, and we leave such a study for future work.
@@ 5 Summary
The lifetime of PBHs is extended by the memory burden effect. Hence, light PBHs may survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs/FIMPs, WIMPs/FIMPs produced via the Hawking radiation of PBHs, and PBHs that survived via the memory burden effect. We focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of PBH particle production via Hawking radiation.
Focusing on parameter regions in which PBHs never dominate the energy density of the Universe, we derived the relic abundance of DM and identified a sufficient condition under which DM particles emitted by PBHs do not enter thermal equilibrium with the thermal bath. In this regime, the contribution from PBH evaporation remains independent of the standard FO (or FI) process, and the total relic abundance of DM can be consistently obtained as a sum of the three components.
In addition, we have examined the potential contribution from gravitational FI via graviton exchange and demonstrated that it remains subdominant within the parameter space considered in this work.
A more detailed numerical investigation of scenarios close to thermalization boundaries or a systematic scan over the full memory burden parameter space would be an interesting direction for future work.